OpenStudy (anonymous):

The distance (in miles) traveled when driving at a certain speed s for 23 hours, then driving miles/hour faster for another hour. Express the distance in terms of s. We'll do it in parts: the distance traveled when driving at s mph for 23 hours is the expression __________ . ADDED to that distance is driving 1 hour at 14 mph FASTER, so the expression to be added is _______. The final answer (expression) then is __________. Please help me on this one!

OpenStudy (anonymous):

d = s x t (speed x time) the distance traveled when driving at s mph for 23 hours is the expression 23s the expression to be added is 1(s+14) so the final answer d = 23(s) + 1(s+14) = 23s + s + 14 so d=24s+14